# Fine-Tuning Masked Diffusion for Provable Self-Correction

### v3

@@ D.1 Omitted Results
Here, we present the comprehensive results on OpenWebText, on MAUVE score, Generative Perplexity (Gen PPL), and Entropy. The graph in Figure 4 is derived from Table 3, PRISM-loop for $N>128$ and PRISM for $N=128$.
Table 3: Evaluation result for PRISM-loop/PRISM/baselines. The best MAUVE scores are bolded. Baseline results marked with are taken from wang2025remasking.
@@
Data 1.00 14.8 5.44
N=1024 N=2048 N=4096 N=1024 N=2048 N=4096 N=1024 N=2048 N=4096
MDLM 0.042 0.037 0.035 51.3 51.3 50.9 5.46 5.46 5.45
ReMDM-conf 0.051 0.060 0.065 48.0 44.5 42.0 5.44 5.42 5.39
ReMDM 0.403 0.610 0.656 28.6 22.8 17.6 5.38 5.30 5.20
[…]
N=128 N=256 N=512 N=128 N=256 N=512 N=128 N=256 N=512
MDLM 0.015 0.023 0.031 61.5 55.8 53.0 5.52 5.49 5.48
ReMDM-conf 0.015 0.022 0.037 74.2 66.0 52.5 5.57 5.54 5.49
ReMDM 0.057 0.216 0.350 42.5 30.5 21.1 5.43 5.34 5.21
[…]
N=16 N=32 N=64 N=16 N=32 N=64 N=16 N=32 N=64
MDLM 0.002 0.006 0.011 230.2 100.5 72.1 5.71 5.60 5.55
ReMDM-conf 0.005 0.006 0.009 244.4 128.4 90.3 5.73 5.66 5.60
ReMDM 0.004 0.005 0.011 257.2 126.4 81.3 5.74 5.65 5.58
[…]
@@ D.2 Details on the remasking Strategy for OpenwebText
To avoid complicated hyperparameter designs, we set these hyperparameters to be a function of $N$. The remasking ratio and the remasking activation step are set to $\eta=\frac{2.56}{N}$ and $l_{on}=\lceil 0.5N\rceil$, respectively. The $N$-step PRISM-loop algorithm comprises an initial $N/2$-step generation for the PRISM and then followed by an $N/2$-step loop with $M=20$.
@@ D.3 Ablation on Fine-tuning Hyperparameters-$(k,n_{y})$
We conducted an ablation study to analyze how the interplay between $k$ and $n_{y}$ affects the generation performance. We tested variants of $(k,n_{y})\in\{(4,8),(8,4),(16,2),(32,1)\}$, thus remaining the total number of token updates per batch constant ($k\times n_{y}=32$) for a fair comparison. All other configurations follow Table  4, except that we set nucleus $p=0.9$ to sample $\mathbf{y}$} from $\mathbf{z}$} (during the PRISM fine-tuning).
@@ D.4 Ablation on Fine-tuning Hyperparameters-Nucleus sampling
We further investigate how the sampling distribution used to generate target sequences $\mathbf{y}$} affects the adapter’s fine-tuning process. Specifically, we analyze the impact of nucleus sampling probability $p$ that we use to sample a token ${\mathbf{y}}^{i}$ from $f_{\theta}^{i}(\cdot\,|\,{\mathbf{z}})$. Our hypothesis is that a less restrictive sampling distribution (large $p$) is more beneficial for training the quality estimator than a shrunk one (small $p$).
@@
D.5 Quantitative analysis of PRISM
@@
D.6 Experiment Configurations
@@
D.7 Additional experimental details on LLaDA
@@
Dataset construction. For PRISM fine-tuning, we adopt the opc-sft-stage-2 (Huang2024OpenCoderTO), comprising $\approx 0.1$M challenging Python coding problems. Each example includes instructions and solutions. Thus, in the masking process, we do not mask instruction tokens.
Training configuration. We use AdamW (loshchilov2017decoupled) (learning rate $1.0\times 10^{-4}$, weight decay $0.01$, warmup ratio $0.05$) and use a cosine schedule with $5$ cycles. Training for $100$ epochs takes approximately $30$ hours on $12$ H100 GPUs.
Inference details. For unmasking, we follow nie2025large and employ semi-autoregressive inference: the sequence is partitioned into multiple blocks, and decoding proceeds from the leftmost block. We set the nucleus parameter to $p=0.0$. These two choices are known to be important for competitive coding performance.

### v4

@@ D.1 Omitted Results
Here, we present the comprehensive results on OpenWebText in terms of MAUVE score, Generative Perplexity (Gen PPL), and Entropy. The graph in Figure 3 is derived from Table 2. We select the subset $\mathcal{S}$ based on confidence scores during fine-tuning.
Table 2: Evaluation result for PRISM/baselines across NN. The best MAUVE scores are bolded. Baseline results marked with are taken from (Wang et al., 2025).
@@
64 128 256 512 1024 64 128 256 512 1024 64 128 256 512 1024
MDLM 0.011 0.015 0.023 0.031 0.042 72.1 61.5 55.8 53.0 51.3 5.55 5.52 5.49 5.48 5.46
ReMDM-conf 0.009 0.015 0.022 0.037 0.051 90.3 74.2 66.0 52.5 48.0 5.60 5.57 5.54 5.49 5.44
ReMDM 0.016 0.057 0.216 0.350 0.403 60.4 42.5 30.5 21.1 28.6 5.51 5.43 5.34 5.21 5.38
PRISM (ours) 0.132 0.278 0.419 0.434 0.510 29.4 23.9 21.6 20.6 20.2 5.37 5.32 5.29 5.27 5.25
@@ D.2 Details on the remasking Strategy for OpenwebText
To avoid complicated hyperparameter designs, we set these hyperparameters to be a function of $N$. The remasking ratio and the remasking activation step are set to $\eta=\frac{2.56}{N}$ and $l_{on}=\lceil 0.5N\rceil$, respectively.
@@ D.3 Ablation on Fine-tuning Hyperparameters-$(k,n_{y})$
We conducted an ablation study to analyze how the interplay between $k$ and $n_{y}$ affects the generation performance. We tested variants of $(k,n_{y})\in\{(4,8),(8,4),(16,2),(32,1)\}$, thus remaining the total number of token updates per batch constant ($k\times n_{y}=32$) for a fair comparison. All other configurations follow Table  5, except that we set nucleus $p=0.9$ to sample $\mathbf{y}$} from $\mathbf{z}$}, and select $\mathcal{S}$ randomly during PRISM fine-tuning.
@@ D.4 Ablation on Fine-tuning Hyperparameters-Nucleus sampling
We further investigate how the sampling distribution used to generate target sequences $\mathbf{y}$} affects the adapter’s fine-tuning process. Specifically, we analyze the impact of nucleus sampling probability $p$ that we use to sample a token ${\mathbf{y}}^{i}$ from $f_{\theta}^{i}(\cdot\,|\,{\mathbf{z}})$, while selecting $\mathcal{S}$ randomly. Our hypothesis is that a less restrictive sampling distribution (large $p$) is more beneficial for training the quality estimator than a shrunk one (small $p$).
@@
D.5 Quantitative analysis of PRISM: Calibration study
@@
D.6 Learning likelihood instead of scalar
Table 3: Comparison between scalar-quality learning and likelihood-learning variants on OpenWebText. PRISM learns a scalar per-token quality score, whereas likelihood-based variants learn the full per-position conditional likelihood. Under the same 10k-step training budget, scalar-quality PRISM is more training-efficient and achieves better MAUVE score. With longer training, likelihood learning becomes competitive, and entropy-aware remasking further improves MAUVE across decoding budgets.
Method MAUVE ($\uparrow$) Gen PPL. ($\downarrow$) Entropy ($\uparrow$)
64 128 256 512 1024 64 128 256 512 1024 64 128 256 512 1024
PRISM (scalar, 10k) 0.132 0.278 0.419 0.434 0.510 29.4 23.9 21.6 20.6 20.2 5.37 5.32 5.29 5.27 5.25
Likelihood (10k) 0.083 0.141 0.179 0.214 0.185 18.8 15.5 14.1 13.5 13.1 5.18 5.13 5.10 5.08 5.07
Likelihood (20k) 0.144 0.249 0.359 0.435 0.421 18.6 15.2 13.7 13.0 12.7 5.20 5.15 5.11 5.09 5.08
Likelihood + Entropy (20k) 0.195 0.322 0.433 0.442 0.520 20.7 18.4 17.5 17.1 16.9 5.28 5.25 5.22 5.21 5.20
Our formulation of PRISM learns a scalar per-token quality score $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g^{i}_{\star}({\mathbf{y}})$}. This scalar is sufficient for identifying low-quality tokens while remaining lightweight and efficient to train. A natural extension is to learn the full per-position likelihood instead:
$p({\mathbf{x}}^{i}=\cdot\mid{\mathbf{y}}\oplus\mathbf{m}_{i})\in\triangle(\mathcal{V}).$
In this section, we study this likelihood-learning variant and analyze the trade-off it introduces for self-correction and remasking. To keep notation consistent with the main text, we slightly overload $g_{\phi}^{i}$: in this subsection, $g_{\phi}^{i}(\cdot\mid{\mathbf{y}}\oplus\mathbf{m}_{i})$ denotes a distribution-valued likelihood head, whereas in the main body $g_{\phi}^{i}({\mathbf{y}})$ denotes a scalar per-token quality score.
Likelihood Learning. The theoretical framework of PRISM extends directly to likelihood learning. Specifically, replacing the BCE objective in Equation 2 with a cross-entropy objective gives
$\mathcal{L}_{CE}(\phi)\coloneqq\mathbb{E}_{{\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i)}\left[-\log{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{x}}^{i}\mid{\mathbf{y}}\oplus\mathbf{m}_{i})}\right],$ (3)
whose minimizer is the conditional likelihood $p({\mathbf{x}}_{i}=\cdot\mid{\mathbf{y}}\oplus\mathbf{m}_{i})$. Unlike PRISM, this variant does not require an additional scalar-quality head. Instead, we can fine-tune the MDM backbone using the existing unmasking head, replacing $g_{\phi}$ in Equation 3 with $f_{\theta}$. Compared to scalar-quality learning, likelihood learning provides richer information at each position, but it is also more expensive to learn as it requires learning a full vocabulary distribution rather than a single scalar.
To study this trade-off, we train likelihood-based variants under the same OpenWebText setup used in our main experiments, but with the likelihood objective above and without an additional scalar-quality head. Table 3 shows that likelihood learning is effective, but underperforms scalar-quality PRISM under the same 10k-step training budget. With longer training, however, the likelihood-based variant becomes competitive with scalar-quality PRISM, indicating that the likelihood objective is compatible with the PRISM framework but requires more optimization.
Figure 7: Relationship between per-token quality and per-position entropy during OpenWebText generation. Each point corresponds to a revealed token during inference. In this setting, remasking is activated only during the 128–256 step interval. Across denoising steps, many tokens exhibit low quality but high entropy, indicating positions where the current token may be plausible despite receiving a low scalar-quality score.
Richer supervision from learned likelihoods.
A key advantage of likelihood learning is that it exposes uncertainty information through the per-position entropy
$H^{i}_{\theta}({\mathbf{x}}_{t}):=-\sum_{v\in\mathcal{V}}g_{\theta}^{i}({\mathbf{x}}_{t})[v]\log g_{\theta}^{i}({\mathbf{x}}_{t})[v],$
This entropy captures the uncertainty of the token distribution at a revealed position $i$. Low per-token quality can arise from two qualitatively different situations: (1) the revealed token is genuinely erroneous, or (2) the token is semantically valid but belongs to a high-entropy position with many plausible alternatives. Remasking tokens in the second case can be redundant or even harmful. Figure 7 shows that low-quality, high-entropy tokens appear throughout the denoising process, suggesting that scalar quality alone does not distinguish these two cases.
Figure 8: Effect of entropy-aware remasking during OpenWebText generation. In this setting, remasking is activated only during the 128–256 step interval. Using the entropy-adjusted score $s^{i}({\mathbf{x}}_{t})$ biases remasking toward low-quality, low-entropy tokens, which are more likely to correspond to genuine errors.
Motivated by this observation, we consider the following entropy-aware remasking score:
$s^{i}({\mathbf{x}}_{t})=g_{\theta}^{i}({\mathbf{x}}_{t})[{\mathbf{x}}_{t}^{i}]\cdot\exp(H^{i}_{\theta}({\mathbf{x}}_{t}))=\exp\left[\log g_{\theta}^{i}({\mathbf{x}}_{t})[{\mathbf{x}}_{t}^{i}]-\mathbb{E}_{v}\left[\log g_{\theta}^{i}({\mathbf{x}}_{t})[v]\right]\right],$
This score compensates for high-entropy positions and therefore avoids unnecessarily remasking tokens that have many plausible alternatives. Using the same remasking schedule with $s^{i}(x_{t})$, we observe that low-quality, low-entropy positions are more selectively remasked, as shown in Figure 8. Quantitatively, the entropy-aware score improves generation quality over the likelihood-only score in Table 3, with especially large gains in the low-sampling-step regime.
That said, this richer supervision comes with a training-efficiency trade-off. On OpenWebText, we observe an approximately $2\times$ training-efficiency gap between scalar-quality learning and likelihood learning. It remains an open question how this gap evolves at larger scales, where learning a full likelihood may become more costly.
D.7 Quantitative analysis of PRISM: Error correction study
Case-by-case analysis.
We provide a qualitative breakdown of PRISM’s self-correction behavior across representative HumanEval tasks. The code snippets are given in Table D.7.
  • Syntax error correction (task id 56): The baseline model produces a syntactically invalid program due to an incomplete elif statement, which results in a SyntaxError. PRISM successfully detects this low-quality token and replaces the invalid branch with a valid else clause, yielding a syntactically correct and functionally valid implementation.
  • Logical error correction (task id 5): The baseline implementation incorrectly appends the delimiter after every element, including the last one. PRISM revises the code to conditionally insert the delimiter only between elements, correcting the logical structure of the program while preserving the intended functionality.
  • (Failure case) overcorrection (task id 25): Although the baseline solution is correct, PRISM rewrites the code into a more complex form that relies on repeated min and max operations. This transformation introduces an error when the input list becomes empty, illustrating a failure mode where PRISM overcorrects a correct but non-canonical solution.
  • (Failure case) global reasoning error (task id 17): Both the baseline model and PRISM fail on this task. The core issue lies in the iteration structure: the program should iterate over multi-character musical symbols rather than individual characters. This type of global structural error is not captured by PRISM’s token-level quality estimation, highlighting a limitation of local self-correction.
Baseline PRISM
Syntax error correction (task 56) def correct_bracketing(brackets): stack = [] for bracket in brackets: if bracket == <: stack.append(bracket) elif: # SyntaxError if not stack: return False stack.pop() return not stack def correct_bracketing(brackets): stack = [] for bracket in brackets: if bracket == <: stack.append(bracket) else: if len(stack) == 0: return False stack.pop() return len(stack) == 0
Logical error correction (task 5) def intersperse(numbers, delimeter): result = [] for num in numbers: result.append(num) result.append(delimeter) return result def intersperse(numbers, delimeter): result = [] for num in numbers: if not result: result.append(num) else: result.append(delimeter) result.append(num) return result
Failure case (task 25) def strange_sort_list(nums): if not nums: return [] nums.sort() result = [] while nums: result.append(nums.pop(0)) if nums: result.append(nums.pop(-1)) return result def strange_sort_list(nums): result = [] while nums: min_val = min(nums) result.append(min_val) nums.remove(min_val) max_val = max(nums) result.append(max_val) nums.remove(max_val) return result
Global reasoning failure (task 17) def parse_music(s): out = [] for note in s: if note == o’: out.append(4) elif note == o|’: out.append(2) elif note == ’.|’: out.append(1) return out def parse_music(s): out = [] for note in s: if note == o’: out.append(4) elif note == o|’: out.append(2) elif note == ’.|’: out.append(1) return out
Table 4: Qualitative comparison of self-correction behavior.
D.8 Experiment Configurations
@@
D.9 Additional experimental details on LLaDA
@@
Dataset construction. For PRISM fine-tuning, we adopt the opc-sft-stage-2 (Huang et al., 2024), comprising $\approx 0.1$M challenging Python coding problems. Each example includes instructions and solutions. Thus, in the masking process, we do not mask instruction tokens.
Training configuration. We use AdamW (Loshchilov and Hutter, 2017) (learning rate $1.0\times 10^{-4}$, weight decay $0.01$, warmup ratio $0.05$) and use a cosine schedule with $5$ cycles. Training for $100$ epochs takes approximately $30$ hours on $12$ H100 GPUs.
Inference details. For unmasking, we follow Nie et al. (2025) and employ semi-autoregressive inference: the sequence is partitioned into multiple blocks, and decoding proceeds from the leftmost block. We set the nucleus parameter to $p=0.0$. These two choices are known to be important for competitive coding performance.